OpenStudy (faolan):

help? fan n medal Find the roots of the equation below. 7x^2 + 1 = 0

OpenStudy (owlcoffee):

In order to find the roots of the equation you show, we have to take in consideration the goal we are aiming for. The roots are indeed, the values of x where the function intersect the x-axis and at the same time y=0. But, that's already covered, we can move on to the equation at hand: \[7x ^2 +1 =0 \] We can just limit ourselves to solving for "x" here, and in order to do that, we apply the reciprocal or direct operations in order to leave "x" alone on either sides of the "=" sign, we begin this by sustracting "1" on both sides: \[7x^2+1-1=0-1 \iff 7x^2=-1\] of course, now we have to get rid of the "7" that multiplies the "x" on the left side, so we can divide both sides by "7": \[\frac{ 7x^2 }{ 7 }=\frac{ (-1) }{ 7 } \iff x^2=-\frac{ 1 }{ 7 }\] The rest is pretty straight forward, but I suggest you take the plus/minus square root of both sides, thereby finding the solution for the equation, and since the whole function relied solely on "x" and it was equal to zero, we can safely say that the solutions for the equation represent the roots of such.

OpenStudy (faolan):

thx